OpenStudy (grant330sims):

Find the polynomial function with roots 1, –2, and 5

OpenStudy (anonymous):

start with the polynomial in factored form, then multiply out

OpenStudy (anonymous):

When x = 1,-2 and 5 and f(x) = 0, then the factors are (x - 1)(x + 2)(x - 5) = 0. (x - 1)(x + 2)(x - 5) = 0 Multiplying each term by term and combining like terms, we obtain: (x² + x - 2)(x - 5) = 0 x³ - 5x² + x² - 5x - 2x + 10 = 0 x³ - 4x² - 7x + 10 = 0 Hence, f(x) = x³ - 4x² - 7x + 10. I hope this helps!

OpenStudy (anonymous):

since you know the zeros are 1, -2 and 5, you know it factors as \[p(x)=(x-1)(x+2)(x-5)\] although you do not know the leading coefficient is 1, so you are finding A polynomial, not THE polynomial